OpenStudy (carissa15):

If anyone could me, I don't know what the mean value theorem is but I have a question about it as below.

OpenStudy (carissa15):

Apply the mean value Theorem to f(x)=In(1+x) to show that \[\frac{ x }{ 1+x }<\ln (1+x)<x, for x>0.\]

OpenStudy (freckles):

\[\text{ try something like this } \\ \text{ use the mvt for } f(x)=\ln(1+x) \text{ on } (0,x)\]

OpenStudy (carissa15):

I am not sure how to use the mean value theorem

OpenStudy (freckles):

the function f(x)=ln(x+1) is continuous on [0,1] and differentiable on (0,1) so there exist \[c \in (0,x) \text{ such that } f'(c)=\frac{f(x)-f(0)}{x-0}\]

OpenStudy (freckles):

and you are definitely going make use that c is in the interval (0,x) that is 0<c<x <--this will be handy inequality in your proof

OpenStudy (freckles):

the function f(x)=ln(x+1) is continuous on [0,x] and differentiable on (0,x)*

OpenStudy (freckles):

anyways let me know if you still need help

OpenStudy (carissa15):

Thank you, so all it is asking is to prove that the function will be continuous and differentiable from the values using the mvt?

OpenStudy (freckles):

no you have to show what it asked

OpenStudy (freckles):

use the thing above find f'(c) and...f(x) and f(0)

OpenStudy (freckles):

\[f(x)=\ln(x+1)\\ \text{ can you find } f'(x)?\]

OpenStudy (freckles):

not sure if you are there or not but I have to go

OpenStudy (danjs):

The mean value theorem basically says, if you have a continuous function between two x values, There is some point in between the interval ends where the slope of the tangent line to the function is the same as the slope of a secant line connecting the endpoints...if i remember right

OpenStudy (danjs):

It is an existence type thing

OpenStudy (danjs):